OpenStudy (anonymous):

How was the derivative for the (circled function) found? Only part i don't understand was it the quotient rule?

OpenStudy (amistre64):

quotient rule is overrated since a quotient is just another type of prduct

OpenStudy (amistre64):

\[\frac fg=fg^{-1}\]

OpenStudy (amistre64):

... also, we have no idea what the (circled function) is to start with :/

OpenStudy (anonymous):

my bad forgot to post it

OpenStudy (amistre64):

thats an awful thing to post ...

OpenStudy (amistre64):

f(g) to f'(g) * g'

OpenStudy (amistre64):

\[f=4u(u^2+3)^{-1}\] \[f'=4(u^2+3)^{-1}-4u(2u)(u^2+3)^{-2}\] \[g=4x^2+7x+1\] \[g'=8x+7\]

OpenStudy (amistre64):

g(0)=1 so define f'(1) and g'(0)

OpenStudy (amistre64):

and yes, they used the "qoutient rule" on your circled part ...

OpenStudy (amistre64):

if you want to use it; i like to define it in terms of t and b, for top and bottom\[\frac tb\to\frac{bt'-b't}{b^2}\]

OpenStudy (anonymous):

i see, lol i understand the quotient rule its just when i tried it i seemd to not get the answer they came out with

OpenStudy (anonymous):

never minddd! lol i found out